# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

### v1

@@ A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing algorithm design in HyGen. For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness. A utility value between 0 and 1 is used to balance the chance between these two data structures. During scheduling, prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, and keeps their original order when scheduling running requests, effectively utilizing prefix sharing while avoiding starvation.
@@
5:     for $r\in R$ do
6:        if $r$.state == DECODE then
7:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
8:           if $t>t_{req}$ then
9:              break
10:           end if
11:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
12:        else if $r$.state == PREFILL then
13:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
14:           if $l>0$ then
15:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
16:           else
17:              break
18:           end if
@@
20:     end for
21:     while $T_{p}$ is not empty do
22:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
23:        if $rand<u$ then
24:           $r\leftarrow T_{p}$.get_next_request()
25:        else
26:           $r\leftarrow T_{f}$.get_next_request()
27:        end if
28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
29:        if $l>0$ then
30:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
31:           if $rand<u$ then
32:              $T_{f}$.remove_request($r$)
33:           else
34:              $T_{p}$.remove_request($r$)
35:           end if
36:        else
37:           break
38:        end if
39:     end while
40:     return $B$
41:  end function

### v2

@@ A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing maximization algorithm design in HyGen. We construct a prefix tree $T_{p}$ to capture prefix sharing charateristics among offline requests. During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled. Additionally, running requests keep their original DFS order in future scheduling process, effectively utilizing prefix sharing.
Algorithm 3 Prefix-sharing-aware offline scheduler
1:  Construct prefix tree $T_{p}$
2:  function PREFIX_SHARING_OFFLINE_SCHEDULE
3:     Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$
4:     Output: batched requests $B$
5:     $B\leftarrow\{\}$
6:     for $r\in R$ do
7:        if $r$.state == DECODE then
8:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:           if $t>t_{req}$ then
10:              break
11:           end if
12:           $t\leftarrow t-t_{req}$
13:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
14:        else if $r$.state == PREFILL then
15:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
16:           if $l>0$ then
17:              $t\leftarrow t-t_{req}$
18:              $c\leftarrow c-l$
19:              $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
20:              $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
21:           else
22:              break
23:           end if
24:        end if
25:     end for
26:     while $T_{p}$ is not empty do
27:        $r\leftarrow T_{p}$.get_next_request()
28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
29:        if $l>0$ then
30:           $t\leftarrow t-t_{req}$
31:           $c\leftarrow c-l$
32:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
33:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
34:           $T_{p}$.remove_request($r$)
35:        else
36:           break
37:        end if
38:     end while
39:     return $B$
40:  end function
A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
This section details the extended version of the SLO-aware prefix sharing maximization algorithm, enhanced with fairness-aware scheduling. For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness. A utility ratio between 0 and 1 is used to balance the chance between these two data structures. During scheduling, the extended prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, based on the utility ratio. The selected request is then deleted from both data structures to ensure synchronization. This solution balances prefix sharing and fairness without disrupting the DFS order of the prefix tree while avoiding possible starvation.
@@ A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
5:     $B\leftarrow\{\}$
6:     for $r\in R$ do
7:        if $r$.state == DECODE then
8:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:           if $t>t_{req}$ then
10:              break
11:           end if
12:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
13:        else if $r$.state == PREFILL then
14:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
15:           if $l>0$ then
16:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
17:           else
18:              break
@@
20:        end if
21:     end for
22:     while $T_{p}$ is not empty do
23:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
24:        if $rand<u$ then
25:           $r\leftarrow T_{p}$.get_next_request()
26:        else
27:           $r\leftarrow T_{f}$.get_next_request()
28:        end if
29:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
30:        if $l>0$ then
31:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
32:           $T_{p}$.remove_request($r$)
33:           $T_{f}$.remove_request($r$)
34:        else
35:           break
36:        end if
37:     end while
38:     return $B$
39:  end function
Appendix B Mooncake Trace Analysis
This section provides an analysis of Mooncake trace, a production trace taken from real-world server with request lengths and request arrival timestamps. Figure 16 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods. This analysis further demonstrates the fluctuant characteristic of LLM services.
(a) 1 hour
(b) 10 minutes
Figure 16: Request rate varies in Moonshot Mooncake’s LLM service over one-hour and ten-minute periods.